A soil analysis was done for the exchangeable cations in the soil and the results were as follows:

Ca = 7 cmol/kg

Mg = 4 cmol/kg

K = 585 mg/kg   (atomic wt. Potassium is 39 g/mol)

Na = 1 cmol/kg

Al= 135 mg/kg   (atomic wt. Aluminum is 27 g/mol)

H= 3 cmol/kg

Calculate the CEC of this soil –do not forget the units and please show example calculations
Answer #1

First we would convert the units to meq /100 gm

Ca = 7 cmol/kg = 7 meq /100 gm

Mg = 4 cmol/kg=4 meq /100 gm

K = 585 mg/kg = (585 / 390) cmol/kg = 1.5 cmol/kg =1.5 meq /100 gm

Na = 1 cmol/kg=1 meq /100 gm

Al= 135 mg/kg = (135 / 270) cmol/kg = 0.5 cmol/kg =0.5 meq /100 gm

H= 3 cmol/kg=3 meq /100 gm

Hence, CEC of the soil is = (7 + 4 + 1.5 + 1 + 0.5 + 3) meq /100 gm = 17 meq /100 gm